# MolFORM: Multi-modal Flow Matching for Structure-Based Drug Design

### v2

@@ 3.2 Multi-modal Flow matching
To promote accurate geometric alignment between predicted and ground-truth molecular structures, we incorporate a Chamfer loss defined over atomic point clouds. Given two point sets $\hat{\mathbf{x}}_{1}=\{\hat{x}_{i}\}_{i=1}^{N}$ and $\mathbf{x}_{1}=\{x_{j}\}_{j=1}^{M}$ representing predicted and reference atomic positions respectively, the Chamfer distance is defined as:
$\mathcal{L}_{\text{Chamfer}}=\$$\frac{1}{N}\sum_{\hat{x}\in\hat{\mathbf{x}}_{1}}\min_{x\in\mathbf{x}_{1}}\|\hat{x}-x\|_{2}$
$+\frac{1}{M}\sum_{x\in\mathbf{x}_{1}}\min_{\hat{x}\in\hat{\mathbf{x}}_{1}}\|x-\hat{x}\|_{2}.$ (8)
This bidirectional distance encourages each predicted atom to be close to some reference atom, and vice versa, effectively guiding the flow matching process to learn a distribution that preserves spatial fidelity to the ground-truth.
The final loss is shown below:
$\mathcal{L}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{type}}+\lambda\cdot\mathcal{L}_{\text{Chamfer}}$ (9)
where $\lambda$ is a weighting hyperparameter that balances the contribution of the Chamfer loss.
@@
Table 1: Combined results for binding affinity, chemical properties, and geometry/clash metrics. (↑)/(↓) denote better. Top 2 results are marked in bold and underlined. The result from baseline model is quoted from  Lin et al. (2024).
3.3 Sampling
We consider two types of generative sampling procedures, corresponding to the discrete atom types and continuous atomic coordinates.
Continuous sampling (atom positions)
For atomic coordinates, we simulate trajectories via the learned velocity field $v_{\theta}(\mathbf{x}_{t},t)$ using Euler integration in $N$ steps:
$\mathbf{x}_{t+\frac{1}{N}}=\mathbf{x}_{t}+\frac{1}{N}\cdot v_{\theta}(\mathbf{x}_{t},t),\quad t\in\left\{0,\tfrac{1}{N},\dots,\tfrac{N-1}{N}\right\}.$ (10)
Starting from $\mathbf{x}_{0}\sim p_{0}$, where $p_{0}$ is a Gaussian distribution, we integrate the dynamics to obtain the final conformation $\mathbf{x}_{1}$.
Discrete sampling (atom types)
Given an initial sample $v_{0}\sim p_{0}$, where $p_{0}$ is a discrete uniform distribution over the number of atom types, we simulate the forward trajectory using Euler integration:
$v_{t+\Delta t}\sim\mathrm{Cat}\left(\delta_{v_{t}}+R_{t}(v_{t},\cdot)\,\Delta t\right).$ (11)
The rate matrix can be computed as an expectation over conditional flows:
$R_{t}(v_{t},j)=\mathbb{E}_{v_{1}\sim q(v_{1}\mid v_{t})}\left[R_{t}(v_{t},j\mid v_{1})\right],$ (12)
where the posterior is given by
$q(v_{1}\mid v_{t})=\frac{\pi_{t}(v_{t}\mid v_{1})\cdot p_{\mathrm{data}}(v_{1})}{p_{t}(v_{t})}.$ (13)
3.4 Direct Preference Optimization(DPO) on Multi-Flow
Building upon these developments, we introduce a Multi-flow DPO approach that jointly optimizes both continuous and discrete data using Reinforcement Learning from Human Feedback (RLHF).
Given winning samples $x_{1}^{w}$ and losing samples $x_{1}^{l}$ for a given sample $x_{1}$ under a preference $\mathcal{P}$, we construct a dataset of preference pairs $\mathcal{D}=\{(p,x_{1}^{w},x_{1}^{l})\}$, where $p$ denotes the protein condition. Using the loss function in Equation (14), we bypass the need to explicitly estimate a reward function and instead directly optimize the parameterized model $p_{\theta}$. Here, $\log\sigma$ denotes the log-sigmoid function, $\beta$ is a scaling (regularization) coefficient, $p_{\theta}$ is the learnable model, and $p_{\text{ref}}$ is a fixed reference model.
$L_{\mathrm{DPO}}(\theta)={}$ $-\mathbb{E}_{(x_{1}^{w},x_{1}^{l})\sim\mathcal{D},{p^{\theta}_{1|t}(x_{1}|x_{t})},t\sim\mathcal{U}[0,1]}\Big[\log\sigma$
$\quad\Big(\beta\log\frac{p_{\theta}(x_{1}^{w}\mid p)}{p_{\text{ref}}(x_{1}^{w}\mid p)}-\beta\log\frac{p_{\theta}(x_{1}^{l}\mid p)}{p_{\text{ref}}(x_{1}^{l}\mid p)}\Big)\Big]$ (14)
The DPO loss can be decomposed into three parts 15, 16 and 17 as below, aligning with Equation (9), to reflect discrete and continuous components:
$L_{\mathrm{DPO}}^{x}(\theta)={}$ $-\mathbb{E}_{(x_{1}^{w},x_{1}^{l})\sim\mathcal{D},\;t\sim\mathcal{U}[0,1],{p^{\theta}_{1|t}(x_{1}|x_{t})}}$
$\quad\Big[\log\sigma\Big(-\beta\big(\|x_{1}^{w}-\hat{x}_{1,\theta}^{w}\|^{2}-\|x_{1}^{w}-\hat{x}_{1,\text{ref}}^{w}\|^{2}$
$\quad-\|x_{1}^{l}-\hat{x}_{1,\theta}^{l}\|^{2}+\|x_{1}^{l}-\hat{x}_{1,\text{ref}}^{l}\|^{2}\big)\Big)\Big]$ (15)
$L_{\mathrm{DPO}}^{\text{point cloud}}(\theta)={}-\mathbb{E}_{(x_{1}^{w},x_{1}^{l})\sim\mathcal{D},\;t\sim\mathcal{U}[0,1],{p^{\theta}_{1|t}(x_{1}|x_{t})}}$
$\quad\Big[\log\sigma\Big(-\beta\big(\mathcal{L}_{\text{Chamfer}}(x_{1}^{w},\hat{x}_{1,\theta}^{w})-\mathcal{L}_{\text{Chamfer}}(x_{1}^{w},\hat{x}_{1,\text{ref}}^{w})$
$\qquad-\mathcal{L}_{\text{Chamfer}}(x_{1}^{l},\hat{x}_{1,\theta}^{l})+\mathcal{L}_{\text{Chamfer}}(x_{1}^{l},\hat{x}_{1,\text{ref}}^{l})\big)\Big)\Big]$ (16)
$L_{\mathrm{DPO}}^{\text{v}}(\theta)={}$ $-\mathbb{E}_{(x_{1}^{w},x_{1}^{l})\sim\mathcal{D},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big($
$\quad-\beta\big(\mathcal{D}^{\theta}_{\text{ref}}(x_{t}^{w}|x_{1}^{w})-\mathcal{D}_{\text{ref}}^{\theta}(x_{t}^{l}|x_{1}^{l})\big)\Big)\Big]$ (17)
with
$\mathcal{D}^{\theta}_{\text{ref}}(x_{t}|x_{1})={}$ $\sum_{j\neq x_{t}}R_{t}^{q}(x_{t},j|x_{1})\log\frac{R_{t}^{\theta}(x_{t},j)}{R_{t}^{\text{ref}}(x_{t},j)}$
$+R_{t}^{\text{ref}}(x_{t},j)-R_{t}^{\theta}(x_{t},j)\>$ (18)
The values $\hat{x}_{1,\theta}^{l}$ and $\hat{x}_{1,\theta}^{w}$ are the predictions from the model $p_{\theta}$, while $\hat{x}_{1,\text{ref}}^{l}$ and $\hat{x}_{1,\text{ref}}^{w}$ are the corresponding predictions from the reference model $p_{\text{ref}}$.
In the discrete space, the rate matrix $R_{t}{(x_{t},x_{t}+\Delta t)}$ is defined as the expectation of the conditional rate matrix as shown in the equation (19).
$R_{t}(x_{t},x_{t+\Delta t})=\mathbb{E}_{p^{\theta}_{1|t}(x_{1}|x_{t})}\left[R^{q}_{t}(x_{t},x_{t+\Delta t}|x_{1})\right]$ (19)
By choosing the uniform noise initialization, $\mathrm{Uniform}([k])$, the unconditional rate matrix becomes equation (20)(Campbell et al., 2024).
$R_{t}(x_{t},x_{t+\Delta t})$ $=\frac{1}{1-t}\,p^{\theta}_{1\mid t}(x_{1}=x_{t+\Delta t}\mid x_{t}),$ (20)
$\text{for }x_{t+\Delta t}\neq x_{t}$
By substituting equation (20) into equation (18) and equation (17), we can get the final loss of DPO of the discrete space in equation (21).
[…]
$\left.\left.\cdot\left(\log\frac{p_{1|t}^{\theta}(x_{1}^{w}\mid x_{t}^{w})}{p_{1|t}^{\text{Ref}}(x_{1}^{w}\mid x_{t}^{w})}-\log\frac{p_{1|t}^{\theta}(x_{1}^{l}\mid x_{t}^{l})}{p_{1|t}^{\text{Ref}}(x_{1}^{l}\mid x_{t}^{l})}\right)\right)\right]$ (21)
Equation (21) is the DPO formulation for discrete flow matching under uniform noising.
3.5 Time Scheduling
Training scheduling
We adopt a $t$-sampling strategy that emphasizes large time steps to improve the accuracy of local structural details. Following the scheduling method from Geffner et al. (2025), we sample time as
$p(t)=0.02\,\mathcal{U}(0,1)+0.98\,\mathcal{B}(1.9,1.0),$ (22)
where $\mathcal{U}$ and $\mathcal{B}$ denote the uniform and Beta distributions, respectively. This schedule reduces weight on early time steps (where global structure is more variable) and concentrates training on later steps to improve geometric fidelity in atom-level predictions.
Sampling scheduling
In conventional flow matching, the sampling trajectory typically uses a uniform discretization of time $t\in[0,1]$ with $N$ equally spaced steps. However, we observe that the model requires finer granularity near $t=1$, where more detailed structure is synthesized. To better capture this behavior and improve generation quality, we adopt a non-uniform time discretization strategy when integrating the ODE using the Euler method. Specifically, we allocate 60 steps in the interval $[0,0.8]$ and 40 steps in $[0.8,1]$, allowing the model to focus more on the final refinement phase.
4 Experiments
4.1 Experiment Setup

### v3

@@ 3.2. Multi-modal Flow Matching
To promote accurate geometric alignment between predicted and ground-truth molecular structures, we incorporate a Chamfer loss defined over atomic point clouds. Given two point sets $\hat{x}_{1}=\{\hat{x}_{i}\}_{i=1}^{N}$ and $x_{1}=\{x_{j}\}_{j=1}^{M}$ representing predicted and reference atomic positions respectively, the Chamfer distance is
(1) $\mathcal{L}_{\mathrm{Chamfer}}=\frac{1}{N}\sum_{\hat{x}\in\hat{x}_{1}}\min_{x\in x_{1}}\|\hat{x}-x\|_{2}+\frac{1}{M}\sum_{x\in x_{1}}\min_{\hat{x}\in\hat{x}_{1}}\|x-\hat{x}\|_{2}.$
Overall objective.
The final pretraining objective for the multimodal flow is
$\mathcal{L}=\mathcal{L}_{\mathrm{pos}}+\mathcal{L}_{\mathrm{type}}+\lambda\cdot\mathcal{L}_{\mathrm{Chamfer}},$
where $\lambda$ is a weighting hyperparameter.
3.3. Sampling
We consider two generative sampling procedures, the continuous atomic coordinates and discrete atom types. Starting from a joint noise prior $(x_{0},v_{0})$ with $x_{0}\sim p_{0}$ (Gaussian) and $v_{0}\sim\mathrm{Uniform}([K])$, evolving both modalities forward from $t=0$ to $t=1$.
Continuous sampling.
For atomic coordinates, we simulate trajectories via the learned velocity field $\mathbf{v}_{\theta}(x_{t},v_{t},p,t)$ using Euler integration in $N$ steps:
(2) $x_{t+\frac{1}{N}}=x_{t}+\frac{1}{N}\cdot\mathbf{v}_{\theta}(x_{t},v_{t},p,t),\qquad t\in\left\{0,\frac{1}{N},\ldots,\frac{N-1}{N}\right\}.$
Integrating this ODE from $x_{0}$ yields the final conformation $x_{1}$.
Discrete sampling.
For atom types, we simulate a CTMC using an Euler discretization. Given the current type $v_{t}$,
$v_{t+\Delta t}\sim\mathrm{Cat}\!\left(\delta_{v_{t}}+R_{\theta,t}(v_{t},\cdot\mid p)\,\Delta t\right),$
where $R_{\theta,t}$ is the model-induced (unconditional) rate matrix and $\delta_{v_{t}}$ denotes the one-hot vector at $v_{t}$. The unconditional rate can be computed as a posterior expectation over the forward conditional rate matrix $R^{q}_{t}(\cdot,\cdot\mid v_{1})$:
$R_{\theta,t}(v_{t},j\mid p)=\mathbb{E}_{v_{1}\sim p_{\theta,1|t}(\cdot\mid v_{t},p,t)}\left[R^{q}_{t}(v_{t},j\mid v_{1})\right].$
The exact Bayesian posterior satisfies
$q(v_{1}\mid v_{t})=\frac{\pi_{t}(v_{t}\mid v_{1})\,p_{\mathrm{data}}(v_{1})}{p_{t}(v_{t})},$
and in practice we approximate $q(v_{1}\mid v_{t})$ using the learned denoising posterior $p_{\theta,1|t}(v_{1}\mid v_{t},p,t)$. (Under the uniform corruption in Eq. (1), this also yields the convenient closed form $R_{\theta,t}(v_{t},j\mid p)=\frac{1}{1-t}\,p_{\theta,1|t}(v_{1}=j\mid v_{t},p,t)$ for $j\neq v_{t}$.
3.4. Direct Preference Optimization (DPO) on Multi-Flow
Multi-flow model provides strong generative capability, downstream objectives (e.g., docking affinity) are not guaranteed to be aligned with the model likelihood. We therefore perform an offline preference-alignment stage using Direct Preference Optimization (DPO). For each protein pocket condition $p$, we collect a winner molecule and a loser molecule and form a preference pair. A molecule at the terminal time is denoted by $m_{1}:=(x_{1},v_{1})$, where $x_{1}$ are continuous atomic coordinates and $v_{1}$ are discrete atom types. The preference dataset is $\mathcal{D}_{\mathrm{pref}}=\{(p,m_{1}^{w},m_{1}^{l})\}$.
Standard DPO objective.
Let $p_{\theta}(m_{1}\mid p)$ be the learnable conditional generative model and $p_{\mathrm{ref}}(m_{1}\mid p)$ be a fixed reference (the pretrained base model). DPO optimizes $\theta$ via a binary classification objective:
$\mathcal{L}_{\mathrm{DPO}}(\theta)$ $=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}}}\Big[\log\sigma\!\Big(\beta\big(\log\tfrac{p_{\theta}(m_{1}^{w}\mid p)}{p_{\mathrm{ref}}(m_{1}^{w}\mid p)}$
$\qquad-\log\tfrac{p_{\theta}(m_{1}^{l}\mid p)}{p_{\mathrm{ref}}(m_{1}^{l}\mid p)}\big)\Big)\Big].$
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the logistic sigmoid and $\beta>0$ controls the strength of preference regularization.
Flow-based surrogate for multimodal data.
For flow matching models, directly evaluating $\log p_{\theta}(m_{1}\mid p)$ is intractable. Following the common practice in diffusion/flow preference alignment, we instead define a timestep-wise surrogate by sampling $t\sim\mathcal{U}[0,1]$ and corrupting each modality with the same forward processes used in pretraining:
$x_{t}\sim q_{t\mid 1}(x_{t}\mid x_{1}),\qquad v_{t}\sim\pi_{t}(v_{t}\mid v_{1}).$
The model predicts the clean sample at time $t$: $\hat{x}_{1,\theta}=\hat{x}_{1,\theta}(x_{t},v_{t},p,t)$ (and analogously $\hat{x}_{1,\mathrm{ref}}$ from the reference model), and predicts a categorical reverse posterior $p_{\theta,1\mid t}(v_{1}\mid v_{t},p)$ (and $p_{\mathrm{ref},1\mid t}$).
To align both continuous coordinates and discrete atom types, we apply DPO separately on each modality and optimize their weighted sum (consistent with the base multi-flow training loss structure). Below we write three modality-wise DPO losses.
DPO on continuous coordinates.
$\mathcal{L}_{\mathrm{DPO}}^{x}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\Delta_{x}^{w}-\Delta_{x}^{l}\big)\Big)\Big],$
where
$\Delta_{x}^{w}$ $=-\big\|x_{1}^{w}-\hat{x}_{1,\theta}^{w}\big\|_{2}^{2}+\big\|x_{1}^{w}-\hat{x}_{1,\mathrm{ref}}^{w}\big\|_{2}^{2},$
$\Delta_{x}^{l}$ $=-\big\|x_{1}^{l}-\hat{x}_{1,\theta}^{l}\big\|_{2}^{2}+\big\|x_{1}^{l}-\hat{x}_{1,\mathrm{ref}}^{l}\big\|_{2}^{2},$
and the noisy states are sampled as $x_{t}^{w}\!\sim q_{t\mid 1}(\cdot\mid x_{1}^{w})$, $x_{t}^{l}\!\sim q_{t\mid 1}(\cdot\mid x_{1}^{l})$, $v_{t}^{w}\!\sim\pi_{t}(\cdot\mid v_{1}^{w})$, $v_{t}^{l}\!\sim\pi_{t}(\cdot\mid v_{1}^{l})$, with $\hat{x}_{1,\theta}^{w}=\hat{x}_{1,\theta}(x_{t}^{w},v_{t}^{w},p,t)$ and $\hat{x}_{1,\theta}^{l}=\hat{x}_{1,\theta}(x_{t}^{l},v_{t}^{l},p,t)$ (and similarly for the reference model).
DPO on point-cloud geometry (Chamfer).
To additionally enforce geometric alignment at the point-cloud level, we apply the same preference objective to the Chamfer distance $L_{\mathrm{Chamfer}}(\cdot,\cdot)$ defined in Eq. 1:
$\mathcal{L}_{\mathrm{DPO}}^{\mathrm{pc}}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\Delta_{\mathrm{pc}}^{w}-\Delta_{\mathrm{pc}}^{l}\big)\Big)\Big],$
where
$\Delta_{\mathrm{pc}}^{w}$ $=-L_{\mathrm{Chamfer}}\!\big(x_{1}^{w},\hat{x}_{1,\theta}^{w}\big)+L_{\mathrm{Chamfer}}\!\big(x_{1}^{w},\hat{x}_{1,\mathrm{ref}}^{w}\big),$
$\Delta_{\mathrm{pc}}^{l}$ $=-L_{\mathrm{Chamfer}}\!\big(x_{1}^{l},\hat{x}_{1,\theta}^{l}\big)+L_{\mathrm{Chamfer}}\!\big(x_{1}^{l},\hat{x}_{1,\mathrm{ref}}^{l}\big).$
DPO on discrete atom types via CTMC rates.
For discrete flow matching, we follow the CTMC rate-matrix parameterization and define a relative rate-based surrogate between the current model and the reference model:
$\mathcal{L}_{\mathrm{DPO}}^{v}(\theta)=-\mathbb{E}_{(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}},\;t\sim\mathcal{U}[0,1]}\Big[\log\sigma\Big(\beta\big(\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}^{w}\mid v_{1}^{w},p,t)\Big$.
$\Big.-\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}^{l}\mid v_{1}^{l},p,t)\big)\Big)\Big].$
The relative rate score $\mathcal{D}^{\theta}_{\mathrm{ref}}$ compares the model-induced unconditional rate matrix $R_{t}^{\theta}(\cdot,\cdot\mid p)$ with the reference rate matrix $R_{t}^{\mathrm{ref}}(\cdot,\cdot\mid p)$ under the forward conditional rate $R_{t}^{q}(\cdot,\cdot\mid v_{1})$:
$\mathcal{D}^{\theta}_{\mathrm{ref}}(v_{t}\mid v_{1},p,t)$ $=\sum_{j\neq v_{t}}\left[R_{t}^{q}(v_{t},j\mid v_{1})\,\log\frac{R_{t}^{\theta}(v_{t},j\mid p)}{R_{t}^{\mathrm{ref}}(v_{t},j\mid p)}\right$.
(3) $\left.\quad+R_{t}^{\mathrm{ref}}(v_{t},j\mid p)-R_{t}^{\theta}(v_{t},j\mid p)\right].$
The model-induced unconditional rate matrix is the posterior expectation of the conditional rate:
$R_{t}^{\theta}(v_{t},j\mid p)=\mathbb{E}_{v_{1}\sim p_{\theta,1\mid t}(\cdot\mid v_{t},p)}\left[R_{t}^{q}(v_{t},j\mid v_{1})\right],\qquad j\neq v_{t},$
and similarly for $R_{t}^{\mathrm{ref}}$ using $p_{\mathrm{ref},1\mid t}$. (As usual in CTMCs, the diagonal is $R_{t}^{\theta}(v_{t},v_{t}\mid p)=-\sum_{j\neq v_{t}}R_{t}^{\theta}(v_{t},j\mid p)$.)
Uniform noising simplification.
Under the uniform corruption initialization $\mathrm{Uniform}([k])$ for the discrete forward process, the unconditional rate admits a closed form:
$R_{t}^{\theta}(v_{t},j\mid p)$ $=\frac{1}{1-t}\,p_{\theta,1\mid t}(v_{1}=j\mid v_{t},p),$
(4) $R_{t}^{\mathrm{ref}}(v_{t},j\mid p)$ $=\frac{1}{1-t}\,p_{\mathrm{ref},1\mid t}(v_{1}=j\mid v_{t},p),\qquad j\neq v_{t}.$
Substituting Eq. (4) into Eq. (3) yields a discrete DPO surrogate expressed purely in terms of the reverse categorical posteriors:
$\mathcal{L}_{\mathrm{DPO}}^{v}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(p,m_{1}^{w},m_{1}^{l})\sim\mathcal{D}_{\mathrm{pref}}\\ t\sim\mathcal{U}[0,1]\end{subarray}}\Bigg[\log\sigma\Bigg(\frac{\beta}{1-t}\Big(\log\frac{p_{\theta,1\mid t}(v_{1}^{w}\mid v_{t}^{w},p)}{p_{\mathrm{ref},1\mid t}(v_{1}^{w}\mid v_{t}^{w},p)}$
$\qquad-\log\frac{p_{\theta,1\mid t}(v_{1}^{l}\mid v_{t}^{l},p)}{p_{\mathrm{ref},1\mid t}(v_{1}^{l}\mid v_{t}^{l},p)}\Big)\Bigg)\Bigg].$
We jointly fine-tune the multi-flow model:
$\mathcal{L}_{\mathrm{MF\text{-}DPO}}(\theta)=\mathcal{L}_{\mathrm{DPO}}^{x}(\theta)+\lambda\,\mathcal{L}_{\mathrm{DPO}}^{\mathrm{pc}}(\theta)+\mathcal{L}_{\mathrm{DPO}}^{v}(\theta),$
where $\lambda$ is the same geometry weighting coefficient.
Figure 2. Online RL on Multi-Flow. By implicitly parameterizing positive and negative branches relative to the reference policy (instead of learning a separate guidance model), we integrate reinforcement guidance directly into the flow-matching objective and steer the generative flow toward high-reward regions.
3.5. Online RL on Multi-Flow
Problem Setup.
In the online reinforcement learning stage, we maintain an anchor (sampling) policy $\pi_{\theta_{\text{old}}}(\cdot\mid p)$ over complete ligand structures $M=(x_{1},v_{1})$ conditioned on a protein pocket $p$. At each iteration, we roll out $K$ candidate molecules $M_{i}\sim\pi_{\theta_{\text{old}}}(\cdot\mid p)$ and evaluate each sample by a biochemical reward $R_{\text{raw}}(M_{i},p)$. Throughout this section, we assume $R_{\text{raw}}$ is higher-is-better; for docking energies such as Vina (lower-is-better), we use a monotone transform (e.g., $\text{Vina}$) as $R_{\text{raw}}$. The complete procedure is summarized in Algorithm 1.
We introduce a latent binary optimality variable $o\in\{0,1\}$, where $o=1$ denotes a high-quality molecule. We define an optimality probability
$r(M,p):=\mathbb{P}(o=1\mid M,p)\in[0,1],$
obtained from a normalized reward transformation (detailed below). The marginal optimality under the anchor policy is
$\mathbb{P}(o=1\mid p)=\mathbb{E}_{M\sim\pi_{\theta_{\text{old}}}(\cdot\mid p)}\big[r(M,p)\big].$
By the law of total probability, the anchor policy admits a mixture decomposition:
$\pi_{\theta_{\text{old}}}(M\mid p)=\mathbb{P}(o=1\mid p)\,\pi^{+}(M\mid p)+\mathbb{P}(o=0\mid p)\,\pi^{-}(M\mid p),$
where $\pi^{+}(M\mid p):=\pi(M\mid o=1,p)$ is the high-reward component and $\pi^{-}(M\mid p):=\pi(M\mid o=0,p)$ is the low-reward component. Our goal is to improve the policy by steering probability mass toward $\pi^{+}(\cdot\mid p)$ while reducing mass on $\pi^{-}(\cdot\mid p)$, without explicit likelihood-ratio estimation.
Reinforcement Guidance Direction.
Following DiffusionNFT, we interpret $r(M,p)$ as a soft indicator that implicitly partitions rollouts from $\pi_{\theta_{\text{old}}}(\cdot\mid p)$ into positive and negative components. Instead of applying policy gradients on intractable likelihoods, DiffusionNFT defines a guided target vector field for the continuous flow:
$v^{*}(x_{t},v_{t},p,t)=v_{\text{old}}(x_{t},v_{t},p,t)+\frac{1}{\beta}\,\Delta(x_{t},v_{t},p,t),$
where $\Delta$ is a reinforcement improvement direction and $\beta>0$ controls the trade-off between staying close to the anchor policy and moving along the improvement direction (the effective guidance strength scales with $1/\beta$). Since $\pi^{+}$ and $\pi^{-}$ are implicit and intractable, we do not explicitly estimate them. Instead, we realize this guidance through an implicit positive/negative branch construction with a fixed mixing coefficient $\beta$, and use $r_{i}$ only for loss reweighting, which yields a stable supervised objective.
Reward Normalization.
For each pocket $p$, we sample $K$ molecules $\{M_{i}\}_{i=1}^{K}$ during rollout. We perform group-centering within each pocket:
$R_{\text{norm}}(M_{i},p):=R_{\text{raw}}(M_{i},p)-\frac{1}{K}\sum_{k=1}^{K}R_{\text{raw}}(M_{k},p).$
We then map the centered reward to an optimality probability $r_{i}\in[0,1]$ via
$r_{i}=0.5+0.5\cdot\mathrm{clip}\!\left(\frac{R_{\text{norm}}(M_{i},p)}{\max(Z,\epsilon)},-1,1\right),$
where $Z>0$ is a normalization scale (e.g., a running estimate of the global reward standard deviation) and $\epsilon$ is a small constant for numerical stability. Group-centering removes pocket-dependent reward bias while preserving within-pocket ranking signals.
Implicit Policy Parameterization.
We optimize on the same forward corruption processes as in pretraining. For each sampled molecule $M_{i}=(x_{1},v_{1})$, we sample $t\sim U[0,1]$ and corrupt both modalities:
$x_{t}\sim q_{t\mid 1}(x_{t}\mid x_{1}),\qquad v_{t}\sim\pi_{t}(v_{t}\mid v_{1}).$
The multi-flow model predicts clean coordinates via the $x$-prediction parameterization and atom-type logits:
$\hat{x}_{1,\theta}(x_{t},v_{t},p,t)$ $=x_{t}+(1-t)\,v_{\theta}(x_{t},v_{t},p,t),$
$\hat{\ell}_{\theta}(x_{t},v_{t},p,t)$ $\in\mathbb{R}^{N_{M}\times K}.$
We compute the corresponding anchor predictions $\hat{x}_{1,\theta_{\text{old}}}$ and $\hat{\ell}_{\theta_{\text{old}}}$ using the frozen anchor network.
Instead of learning separate guidance models, we implicitly construct positive/negative branches relative to the anchor policy. For continuous coordinates, define
$\hat{x}^{+}_{1,\theta}(x_{t},v_{t},p,t)=(1-\beta)\,\hat{x}_{1,\theta_{\text{old}}}(x_{t},v_{t},p,t)+\beta\,\hat{x}_{1,\theta}(x_{t},v_{t},p,t),$
$\hat{x}^{-}_{1,\theta}(x_{t},v_{t},p,t)=(1+\beta)\,\hat{x}_{1,\theta_{\text{old}}}(x_{t},v_{t},p,t)-\beta\,\hat{x}_{1,\theta}(x_{t},v_{t},p,t).$
For discrete atom types, linear interpolation in probability space is invalid, so we apply the same construction in logit (pre-softmax) space:
$\hat{\ell}^{\pm}_{\theta}(x_{t},v_{t},p,t)=(1\mp\beta)\,\hat{\ell}_{\theta_{\text{old}}}(x_{t},v_{t},p,t)\pm\beta\,\hat{\ell}_{\theta}(x_{t},v_{t},p,t).$
Intuitively, matching the positive branch to the ground truth pulls the model toward high-reward directions, while matching the negative branch encourages moving away from low-reward regions. This realizes negative-aware policy improvement without explicitly estimating $\pi^{+}$, $\pi^{-}$, or $\Delta$.
Joint Optimization Objective.
We optimize a reward-weighted supervised objective over the implicit positive and negative branches. Given $(M_{i},p)$, $t$, and $r_{i}$, we combine the continuous position loss $L_{\text{pos}}$ and the discrete cross-entropy loss $L_{\text{CE}}$:
(5) $L_{\text{NFT}}(\theta)=\mathbb{E}_{t,\,M_{i}}\Big[r_{i}\,\ell^{+}_{\text{pos}}+(1-r_{i})\,\ell^{-}_{\text{pos}}+r_{i}\,\ell^{+}_{\text{ce}}+(1-r_{i})\,\ell^{-}_{\text{ce}}\Big],$
where
$\ell^{\pm}_{\text{pos}}=L_{\text{pos}}\!\big(\hat{x}^{\pm}_{1,\theta}(x_{t},v_{t},p,t),\,x_{1}\big),\qquad\ell^{\pm}_{\text{ce}}=L_{\text{CE}}\!\big(\hat{\ell}^{\pm}_{\theta}(x_{t},v_{t},p,t),\,v_{1}\big).$
This objective remains fully supervised on the forward process while enabling policy improvement via reward-based reweighting, avoiding explicit likelihood-ratio estimation or policy-gradient updates.
3.6. Confidence Head
To explicitly estimate the quality of the generated molecules, we design an auxiliary Confidence Head module like PocketXMol (Peng et al., 2024). This optional module consists of two lightweight Multi-Layer Perceptrons (MLPs) that take the final invariant node embeddings from the backbone network as input. Specifically, one MLP predicts the confidence of atom types, formulated as a binary classification task to determine whether the predicted atom type matches the ground truth. The other MLP estimates structural reliability by regressing the spatial deviation between the generated and ground-truth coordinates. During the training phase, the confidence loss is integrated into the total objective as an auxiliary term, allowing the confidence head to be optimized jointly with the generative flow matching task.
4. Experiments
@@
Table 1. Combined results for binding affinity, chemical properties, and geometry/clash metrics. (↑)/(↓) denote better. Top 2 results are marked in bold and underlined. The result from baseline model is quoted from  Lin et al. (2024).
4.1. Experiment Setup